OpenStudy (jessiegonzales):

MEDAL!!!! 7. Solve the absolute value equation. Graph the solution. 2 l3x-5l -8=8

OpenStudy (jessiegonzales):

A. x=-1 5/6 or x=5/6 B. x=-1 or x=4 1/3 C. x=-1 5/6 or x=4 1/3 D. x=-1 5/6 or x=-1

OpenStudy (jessiegonzales):

@nikato

OpenStudy (nikato):

1. add 8 to both sides. 2. divide both sides by 2. now what do you have?

OpenStudy (jessiegonzales):

10 l 4

OpenStudy (jessiegonzales):

2 l3x-5l -8=8 +8 +8 10 l3x-5l 8 /2 /2 5 l 4

OpenStudy (nikato):

 ok?